# Multi-Fidelity Bayesian Optimization With Across-Task Transferable Max-Value Entropy Search

### v3

@@ III-A Multi-Fidelity Max-Value Entropy Search
$\quad-\mathbbm{E}_{p(f_{n}^{*}|\mathbf{x},\bm{\theta},\mathcal{D}_{n,t})}[H(y_{n}^{(m)}|f_{n}^{(m)}(\mathbf{x})\leq f_{n}^{*},\mathbf{x},\bm{\theta},\mathcal{D}_{n,t})],$ (21)
@@ V-A Two-Phase MFT-MES
To facilitate the analysis of the proposed MFT-MES scheme, we study a generally less efficient variant that retains the essential features of the approach. Fixing the fidelity level $M$, the simplified MFT-MES protocol under study splits the iterations for each task $n$ into two separate phases, one devoted to per-task exploration via MES, and the other dedicated to the acquisition of information about the parameter vector $\bm{\theta}^{*}$. Note that MFT-MES in Algorithm 2 adopts a more flexible strategy that integrates these two phases.
@@ VII Conclusion
Future work may address theoretical performance guarantees in the forms of regret bounds to explain aspects such as the dependence of the optimal weight parameter $\beta$ used by MFT-MES as a function of the number of tasks and total query cost budget. Furthermore, it would be interesting to investigate the extensions to multi-objective multi-fidelity optimization problems [55]; the scalability to higher dimensions of the search space [56]; and the potential gains from incorporating generative models [57].

### v4

@@ III-A Multi-Fidelity Max-Value Entropy Search
$\quad-\mathbbm{E}_{p(f_{n}^{*}|\bm{\theta},\mathcal{D}_{n,t})}[H(y_{n}^{(m)}|f_{n}^{(m)}(\mathbf{x})\leq f_{n}^{*},\mathbf{x},\bm{\theta},\mathcal{D}_{n,t})],$ (21)
@@ V-A Two-Phase MFT-MES
To facilitate the analysis of the proposed MFT-MES scheme, we study a generally less efficient variant that retains the essential features of the approach. Fixing the fidelity level $M$, the simplified MFT-MES protocol under study splits the iterations for each task $n$ into two separate phases, one devoted to per-task exploration via MES, and the other dedicated to the acquisition of information about the parameter vector $\bm{\theta}^{*}$. Note that MFT-MES in Algorithm 2 adopts a more flexible strategy that integrates these two phases. Appendix -B elaborates on the impact of the assumed simplifications.
@@ VII Conclusion
Future work may address theoretical performance guarantees in the forms of regret bounds to explain aspects such as the dependence of the optimal weight parameter $\beta$ used by MFT-MES as a function of the number of tasks and total query cost budget. Furthermore, it would be interesting to investigate the extensions to multi-objective multi-fidelity optimization problems [55] or safe optimization problems with unknown constraints [56]; the scalability to higher dimensions of the search space [57]; and the potential gains from incorporating generative models [58].
-A Proof of Theorem 1
We start the proof of Theorem 1 by deriving an upper bound over the discrepancy in the normalized expected regret with known parameter vector $\bm{\theta}^{*}$. Recall that the first set of iterations is devoted for the current task, i.e., the optimizer selects $\mathbf{x}_{n,t+1}=\arg\max_{\mathbf{x}}\mathbb{E}_{\bm{\theta}\sim p(\bm{\theta}|\mathcal{D}_{n-1},\mathcal{D}_{n,t})}[\bar{\alpha}_{\bm{\theta}}^{\text{MES}}(\mathbf{x})]$ for $t+1\leq c\cdot T$. Under Assumption 1, i.e., $\tilde{r}_{n,t}(\mathbf{x}|\bm{\theta})$ is $\epsilon_{1}$-sharp as per Definition 1 as well as $\epsilon_{2}$-stable as per Definition 2, and Assumption 2, we have
$\big{|}\big{|}\mathbf{x}_{n,t}-\arg\min_{\mathbf{x}\in\mathcal{X}}\tilde{r}_{n,t}(\mathbf{x}|\bm{\theta}^{*})\big{|}\big{|}$
$\leq(\varepsilon_{1}+\varepsilon_{2})\frac{\xi}{(n-1)\cdot(1-c)^{\alpha}\cdot T^{\alpha}},\text{ for sufficiently large $n$}$ (55)
from the triangle inequality. Furthermore, by assuming the $L$-Lipschitz continuity of $\tilde{r}_{n,t}(\mathbf{x}|\bm{\theta}^{*})$, we have
$\Delta_{n,t}$ $=\tilde{r}_{n,t}(\mathbf{x}_{n,t}|\bm{\theta}^{*})-\min_{\mathbf{x}\in\mathcal{X}}\tilde{r}_{n,t}(\mathbf{x}|\bm{\theta}^{*})$
$\leq L\cdot\big{|}\big{|}\mathbf{x}_{n,t}-\arg\min_{\mathbf{x}}\tilde{r}_{n,t}(\mathbf{x}|\bm{\theta}^{*})\big{|}\big{|}$
$\leq\frac{\varepsilon\cdot L\cdot\xi}{(n-1)\cdot(1-c)^{\alpha}\cdot T^{\alpha}},$ (56)
where we have defined $\varepsilon=\varepsilon_{1}+\varepsilon_{2}$.
The expected regret of round $t$ at task $n$ is
$\bar{r}_{n,t}(\mathbf{x}_{n,t}|\bm{\theta}^{*})$ $=\hat{m}_{n,t}(\bm{\theta}^{*})-\mu_{\bm{\theta}^{*}}(\mathbf{x}_{t})$
$=\frac{\hat{m}_{n,t}(\bm{\theta}^{*})-\mu_{\bm{\theta}^{*}}(\mathbf{x}_{t})}{\sigma_{\bm{\theta}^{*}}(\mathbf{x}_{t})}\cdot\sigma_{\bm{\theta}^{*}}(\mathbf{x}_{t})$
$=\tilde{r}_{n,t}(\mathbf{x}_{t}|\bm{\theta}^{*})\cdot\sigma_{\bm{\theta}^{*}}(\mathbf{x}_{t})$
$=(\nu_{n,t}+\Delta_{n,t})\cdot\sigma_{\bm{\theta}^{*}}(\mathbf{x}_{t}),$ (57)
where the last equality is obtained from (-A).
Therefore, based on (-A), we have
$\mathcal{R}_{n,cT}(\mathbf{X}_{n,cT}|\bm{\theta}^{*})$ $=\sum_{t=1}^{c\cdot T}(\nu_{n,t}+\Delta_{n,t})\cdot\sigma_{\bm{\theta}^{*}}(\mathbf{x}_{t})$
$\leq(\nu_{n,t_{n}^{*}}+\Delta_{n,t_{n}^{*}})\sum_{t=1}^{c\cdot T}\sigma_{\bm{\theta}^{*}}(\mathbf{x}_{t})$
$\leq(\nu_{n,t_{n}^{*}}+\Delta_{n,t_{n}^{*}})\sqrt{c\cdot T\sum_{t=1}^{c\cdot T}\sigma_{\bm{\theta}^{*}}^{2}(\mathbf{x}_{t})}$
$\leq(\nu_{n,t_{n}^{*}}+\Delta_{n,t_{n}^{*}})\sqrt{\frac{2c\cdot T\cdot\rho_{n,cT}}{\log(1+\sigma^{-2})}},$ (58)
where the second inequality is obtained directly from the definition of $t_{n}^{*}$; the third inequality comes from Cauchy-Schwarz inequality; and the last inequality is a direct consequence of [17, Lemma 5.3] and [16, proof of Theorem 3.2].
-B Discussion on Theorem 1
The theoretical analysis presented in Sec. V applies the following simplifications to MFT-MES:
  1. 1.
    Separate per-task optimization and acquisition of shared knowledge: Each optimization step focuses either on enhancing per-task performance or on reducing the uncertainty on the shared parameter vector $\bm{\theta}$.
  2. 2.
    Myopic shared knowledge acquisition: The acquisition function used in the steps devoted to the reduction of uncertainty on the parameter vector $\bm{\theta}$ relies on a batch-wise mutual information criterion (43) that disregards observations from previous tasks.
  3. 3.
    Two-phase protocol: The iterations are divided into two separate phases, with the first fraction $c$ of iterations used for per-task optimization, and the remaining fraction $1-c$ of iterations devoted to acquire shared knowledge.
The performance of the resulting simplified scheme is expected to be worse than the proposed MFT-MES algorithm, thus intuitively providing an upper bound on the regret performance of MFT-MES.
In order to gain insights into the gap between the performance of the analyzed simplified scheme and of MFT-MES, this appendix evaluates the performance of the following successive simplifications of MFT-MES through a numerical experiment:
  1. 1.
    Separation-based MFT-MES: This scheme applies only the first simplification, namely separate per-task optimization and acquisition of shared knowledge. Accordingly, the next candidate solution is selected as
    $\mathbf{x}_{n,t+1}=\begin{cases}\arg\max\limits_{\mathbf{x}\in\mathcal{X}}\mathbbm{E}_{p(\bm{\theta}|\mathcal{D}_{n-1},\mathcal{D}_{n,t})}[\alpha_{\bm{\theta}}^{\text{MF-MES}}(\mathbf{x},M)]&\\ \quad\text{with probability $c$},\\ \arg\max\limits_{\mathbf{x}\in\mathcal{X}}\frac{\Delta\alpha^{\text{MFT-MES}}(\mathbf{x},M)}{\lambda^{(M)}}&\\ \quad\text{with probability $1-c$}.\end{cases}$ (59)
    Accordingly, the algorithm opts independently at each iteration to either enhance the current task performance or acquire transferable knowledge.
  2. 2.
    Separation-based Myopic MFT-MES: This scheme applies the first two simplifications, selecting the next candidate solution as
[…]
    This implies that the algorithm opts independently at each iteration to either focus on current task optimization or to acquire shared knowledge by adopting the batch-wise mutual information criterion (43).
  3. 3.
    Two-phase MFT-MES (Two-phase): This scheme applies all the simplifications, corresponding to the MFT-MES variant studied in Sec. V-A for theoretical analysis.
Figure 7: Synthetic optimization tasks: Simple regret (53) against the number of tasks, $n$, for MF-MES (black dot-dashed line), Continual MF-MES (blue dashed line), MFT-MES with $\beta=1.5$ (purple solid line), along with three simplified versions of MFT-MES (see Appendix B): Separation-based MFT-MES (red dashed line), Separation-based Myopic MFT-MES (green dashed line) and Two-phase MFT-MES (orange dashed line). The number of SVGD particles is set to $V=10$.
Fig. 7 shows the simple regret (53) as a function of the number of tasks $n$ observed so far. The weight parameter in (37) for MFT-MES is set to $\beta=1.5$, the fraction parameter for the simplified schemes is set to $c=0.6$, and other numerical settings are kept the same as in Sec. VI-B. As illustrated in Fig. 7, each successive simplification deteriorates the performance of MFT-MES. However, the performance gaps are fairly marginal. Furthermore, all the simplified variants of MFT-MES always outperform the benchmarks MF-MES and Continual MF-MES, as long as the number of tasks is larger than $n=2$. This result provides empirical evidence that our theoretical analysis offers relevant insights into the benefits of MFT-MES by studying a simplified scheme that retains the main features of MFT-MES. Further experimental results can be found in the supplementary material.